# Automated Cell Structure Extraction for 3D Electron Microscopy by Deep Learning

### v1

@@ Abstract
Modeling the 3D structures of cells and tissues is crucial in biology. Sequential cross-sectional images from electron microscopy provide high-resolution intracellular structure information. Segmentation of complex cell structures remains a laborious manual task for experts, demanding time and effort. This bottleneck in analyzing biological images requires efficient and automated solutions.
@@
This study explores deep learning-based automated segmentation of biological images, enabling accurate reconstruction of the 3D structures of cells and organelles. We constructed an analysis system for the cell images of Cyanidioschyzon merolae, a primitive unicellular red algae. This system utilizes sequential cross-sectional images captured by Focused Ion Beam Scanning Electron Microscopes (FIB-SEM). We adopted the U-Net and performed pre-training to identify and segment cell organelles from single-cell images. In addition, we employed the Segmentation Anything Model (SAM) and the 3D watershed algorithm to extract individual 3D images of each cell from large-scale microscope images containing numerous cells. Finally, we applied the pre-trained U-Net to segment each structure within these 3D images. Through this procedure, we could fully automate the creation of 3D cell models.
Our approach would apply to other cell types, and we aim to build a versatile analysis system. We will also explore adopting other deep learning techniques and combinations of image processing methods to further enhance segmentation accuracy.
Keywords Deep learning  $\cdot$ FIB-SEM  $\cdot$ U-Net  $\cdot$ Segmentation Anything Model  $\cdot$ 3D images
1 Introduction
Advanced deep learning technology has significantly improved the accuracy and efficiency of general image recognition tasks. However, its application for biological tissue and medical imaging still needs to be improved.
@@ 1 Introduction
As a demonstration, we analyzed 3D cell images captured using a Focused Ion Beam Scanning Electron Microscope (FIB-SEM). FIB-SEM is primarily used to measure physical properties and evaluate microstructures pmid11207928 of solid samples. Recently, it has also been utilized to analyze 3D structures of cells and tissues. The surface of resin-embedded biological samples is repeatedly cut using a gallium ion beam and imaged by Scanning Electron Microscopes. The serial block-face (SBF) image data, i.e., a sequence of cross-sectional images, is reconstructed as a 3D structure.
@@
However, the intricate boundaries separating structures within each slice, low contrast, and significant noise make it challenging to analyze based on brightness. Furthermore, since images taken with an electron microscope are single channel (grayscale), relying on electron beam reflections, the absence of color information further complicates the analysis. In addition, it is also difficult to identify specific cells within a space containing multiple cells and extract their microstructures. As a result, the segmentation of microstructure depends on experts’ subjective evaluation and manual techniques, which consume a lot of time. Image data analysis, which automates this process and is more efficient, objective, and quantitative, is expected to contribute to analyzing various 3D biological images. This study aims to fully automate this process, exploring the application of deep learning in the recognition tasks of cultured cell images.
@@
We chose Cyanidioschyzon merolae as a model for the automated reconstruction of 3D cell structures using FIB-SEM jin_sci ; Takasaka2023 . C. merolae is a primitive unicellular eukaryote belonging to the red algae, with a diameter of approximately 1.5-2.0 $\mu$ m 1010000781756803720 (Figure 1). This organism has a simple structure with only one organelle of each type: the nucleus, mitochondrion, plastid, and peroxisome pmid33836072 . Each organelle is distributed equally to the two daughter cells pmid9522454 ; pmid15681662 . The genome of C. merolae and the genomes of its mitochondrion and plastid have been completely sequenced pmid9801318 ; pmid12755171 ; pmid15071595 ; pmid17623057 . Moreover, it is relatively easy to synchronize the cell division cycles of multiple C. merolae cells in culture with the cycle of light pmid8082652 . Due to these characteristics, C. merolae is regarded as a model organism for investigating the mechanisms of cell division. Cell division in C. merolae has been extensively studied using molecular and cell biological approaches. Recent advances in observation techniques have made it possible to observe dynamic changes in each living cell’s behavior and target molecules pmid33580410 . However, the physical mechanism of division control based on mechanical properties remains unclear.
@@
Therefore, in this study, we propose an automated method for accurately extracting cell and organelle structures during the division process of C. merolae. In previous studies, a method for segmenting each structure from C. merolae FIB-SEM images using machine learning has already been proposed pmid32060319 . However, this method segments structures from each cross-sectional image in the SBF image containing only one cell. Separating cells and backgrounds is still performed manually and is time-consuming. Additionally, the method disregards the sequential contours between cross-section images and has problems in the segmentation of structures at the cell edge. This study proposes a technique to resolve such shortcomings. By using this method, which allows for the automated and rapid acquisition of 3D structures of multiple C. merolae cells, we expect to contribute to investigating mechanical characteristics related to cell division in the future.
(a) SBF image containing one cell
(b) 3D model of a cell
(c) 3D model of organelles
(d) Sequential cross-sectional images of C. merolae
Figure 1: Procedure to reconstruct a 3D model of a cell from sequential cross-sectional images. (a) An SBF image that centrally includes the cell of interest, cropped from FIB-SEM images. (b) By reconstructing this data and removing the background, the 3D model of the cell was produced. (c) The 3D organelle model was visualized by further removing the cytoplasm. Indicating cytoplasm (white), plastid (green), mitochondrion (yellow), nucleus (gray), and peroxisome (orange). (d) This SBF consists of sequential cross-sectional images, as depicted in the top row. Various organelles, the cytoplasm, and the background in these images were manually identified by experts, as shown in the bottom row.
2 Pre-training for Cell Structure Extraction
2.1 Method
2.1.1 Image data acquisition
@@
The images of freeze-fixed samples containing multiple C. merolae cells were obtained by FIB-SEM. The obtained sequential sectional images’ aspect ratios were corrected, and the noise was removed using a median filter using Amira software (Thermo Fisher Scientific Inc.). Experts cropped to construct learning data images to create an SBF image containing one cell with sides approximately 200-500 pixels long, like Figure 1-(a). They manually segmented each sequential cross-sectional image constituting the SBF image to label plastid, mitochondrion, peroxisome, nuclei, cytoplasm, and background 1360004232021292800 . In other words, all voxels constituting the SBF images were labeled into six classes.
@@
The cell cycle of C. merolae is classified into five stages pmid8082652 , shown as Figure 2. The G1-S phase of the cell cycle is referred to as Stage 1, followed by the G2 phase to the M phase (pre-phase) as Stage 2, the M phase (pre-phase) to the M phase (post-phase) as Stage 3, the M phase (post-phase) as Stage 4, and the M phase (termination phase) as Stage 5. Multiple SBF images were labeled for each stage. The labeled SBF images total 68 and are classified into one of these stages by experts.
(a) Stage1: G1-S phase
(b) Stage2: G2 to M pre-phase
(c) Stage3: M pre-phase to M post-phase
(d) Stage4: M post-phase
(e) Stage5: M termination phase
Figure 2: Structural changes associated with the cell cycle of C. merolae. Each structure indicates cytoplasm (white), plastid (green), mitochondrion (yellow), nucleus (gray), and peroxisome (orange). Each type of organelle is equally distributed to the daughter cells during division pmid9522454 ; pmid15681662 . The division of organelle structures occurs at nearly consistent timings regardless of the individual, although the timing varies depending on the type of organelle pmid33580410 .
2.1.2 Image data preprocessing
@@
The images obtained by FIB-SEM and the labeled dataset created by experts were preprocessed for deep learning analysis; due to a difference in resolution between the horizontal and vertical directions when capturing images with FIB-SEM, the voxels in the SBF images are shaped as rectangular parallelepipeds rather than cubes. The SBF images were resized using nearest-neighbor interpolation to restore the aspect ratio of the voxels to their actual size ratio, facilitating 3D data analysis. A bilateral filter technique was employed to process the images and remove noise that Amira software could not eliminate BCCS12 . Each sectional image was resized to create cubic voxels, each side measuring 256 pixels, to ensure a consistent input image size for the deep learning model. For effective deep learning, the brightness values of the electron microscope images were normalized to a range of 0 to 1 by dividing by 255. After these preprocessing steps, the dataset was split into different sets for cross-validation, resulting in six cases. In each case, an SBF image was selected as validation data for pre-training the deep learning model, each corresponding to a different stage of cell division. The remaining SBF images were used as training data for the deep-learning model. In addition to the conventional cross-sectional images obtained from the top surface by FIB shaving, the SBF training image dataset was augmented by creating new cross-sectional images from both lateral and anterior views. Details on the number of cells used and the number of cross-sectional images for each dataset in each cross-validation case are presented in Table 1.
2.1.3 Deep neural network architecture
@@
We developed a scanning U-Net neural network model, as illustrated in Figure 3 NIPS2012_c399862d ; ioffe2015batch . Initially created for image segmentation tasks, U-Net is now extensively used in various fields, including medical imaging 10.1007/978-3-319-24574-4_28 . The first half of the neural network consists entirely of convolutional layers. On the opposite side, deconvolutional layers gradually restore the image size to the original input. The convolutional and deconvolutional layers are connected by skip connections, designed to prevent information loss due to the encoding of the convolutional layers. In this study, we implemented four stages of convolution and deconvolution, with each deconvolution layer in our model enhanced with an attention gate mechanism vaswani2023attention . The ReLU function was used as the activation function at each layer, and batch normalization was performed at all layers to prevent overfitting NIPS2012_c399862d ; ioffe2015batch . The U-Net model in this study had approximately one hundred thirty million parameters. We carried out a six-class segmentation task. The input data shape was defined as a tensor of (batch size, number of channels, vertical pixel count of the SBF image, and horizontal pixel count of the SBF image). The label data and the U-Net output values were fed into the TverskyLoss function for loss score calculation DBLP:journals/corr/abs-1810-07842 . The AdamW learning function was used, with the learning rate starting from a base rate of 0.00001 and decaying by a factor of 0.95 each epoch DBLP:journals/corr/abs-1711-05101 . The number of epochs was set to 5. All networks were built on a GEFORCE RTX 4090 with 24 GB memory, and the framework was implemented using Python’s TensorFlow. To accelerate learning, the GPU’s tensor cores were automatically utilized.
2.1.4 Segmentation accuracy
@@
During the pre-training phase, we evaluated the accuracy of the learning model through cross-validation. In our cross-validation process, we had six distinct cases (Table 1); for each case, a dataset of 63 SBF images was used for training, and a separate set of 5 SBF images (one for each stage in the cell cycle) was used for validation. The validation SBF images did not overlap between cases. Each validation set included one cell from each of the five stages of the cell division cycle. We segmented the sequential cross-sectional images into six classes using the U-Net for each case. The U-Net’s output approached 1 if the segmentation result indicated a specific class and approached 0 otherwise. Finally, we compared these inference results with the label data and evaluated the model using the Intersection over Union (IoU) as Formula (1). The Area of Overlap refers to the intersection area of the predicted and ground truth area. In contrast, the Area of Union refers to the predicted and ground truth together minus the area of overlap.
$IoU=\frac{\text{Area of Overlap}}{\text{Area of Union}}$ (1)
2.2 Result and discussion
Table 3 presents the IoU score for each case based on the U-Net’s inference results. The table organizes six training classes by row and five cell division stages by column. These results indicate that the average accuracy of background identification is approximately 0.99, suggesting that cells and backgrounds can be identified with very high accuracy, as illustrated in Figure 4. Additionally, the accuracy of plastid inference remains high throughout all stages of cell division. In contrast, mitochondrion and nuclei are smaller in volume than plastid, thus posing a more significant challenge for accurate identification, with an average accuracy of about 0.70. The average accuracy for peroxisomes is notably lower, around 0.45. Furthermore, the inference accuracy for all classes tends to decrease gradually as the cell division stage progresses.
Plastids are more easily identifiable due to their size, unique structure, and distinct brightness values. In contrast, mitochondria and nuclei have brightness values similar to those of background and starch, making them more challenging to identify. Additionally, peroxisomes are assumed to be difficult for deep learning models to learn due to their irregular shapes that vary from cell to cell and brightness values similar to background and starch miyagishima1999 ; ijms21155452 . Furthermore, as cell division progresses, especially before division, a single SBF image may encompass two cells, increasing the complexity of cell structures within SBF images and potentially decreasing the inference accuracy as cell division advances.
Figure 4: The example illustrating the observation of a 3D cell model from three directions. Inference results for the validation dataset using a model obtained through pre-training. Each structure indicates cytoplasm (white), plastid (green), mitochondrion (yellow), nucleus (gray), and peroxisome (orange). The output images of the plastid and cytoplasm closely match the ground truth.
2.3 Enhancing Segmentation Accuracy
We segmented various cellular organelles from cross-sectional images of 2D pictures. However, a problem was identified with low identification accuracy in some classes. To address this issue, we generated new cross-sectional images perpendicular to the original images, using the 3D SBF data, as shown in figure 5. We used these images for inference with a pre-trained model. This model calculates the probability of belonging to all classes for each voxel (the smallest unit of a 3D image). Three predictive values are obtained for each voxel with inference from three directions. The class with the highest probability is selected when different classes are predicted. This process may introduce noise, so a median filter removes noise after selection. As a result, as shown in table 3, some cases showed no change or a slight decrease in predictive accuracy, but other samples showed a clear improvement in accuracy. This suggests that it might not be possible to accurately determine the class from a cross-sectional image from a specific direction. Also, considering that the orientation of the cell to be imaged cannot be predetermined and that inference from difficult angles is inevitable, inferring from three directions is worthwhile, even if it means accepting a slight decrease in accuracy.
(a) 3D Model - Captured Direction
(b) Cross-section 1 - Captured image and Label Data
(c) Cross-section 2 - Captured image and Label Data
(d) Cross-section 3 - Captured image and Label Data
Figure 5: Cross-sections and label data of cells in various directions. (a) a 3D cell model reconstructed from captured consecutive cross-sectional images. This model is visualized from the captured direction. (b) includes consecutive cross-sectional images taken near the center of the cell (left) and label data created by experts based on them (right). (c) and (d) display new cross-sectional images generated from stacked consecutive cross-sectional images (left) and the corresponding label data (right). These images were generated from two directions, one aligned with the captured direction and the other perpendicular. Each panel indicates cytoplasm (white), plastid (green), mitochondrion (yellow), nucleus (gray), and peroxisome (orange).
3 Large-Scale Microscopic Image segmentation
3.1 Segmentation Anything Model
@@
In this study, our developed U-Net model utilized sectional images containing a single cell extracted by experts as training data. However, this model is not ideally suited for analyzing images with multiple cells. Therefore, to automate the entire process, it is necessary to develop another method capable of automatically creating images comparable to expert-generated SBFs. To generate SBFs automatically, we initially utilized SAM to distinguish cells and backgrounds automatically. Developed by Meta, the Segmentation Anything Model (SAM) enables easy segmentation of cells and backgrounds from large-scale microscopic images containing multiple cells kirillov2023segment . SAM has been trained on over 110 million images using over one billion masks and possesses zero-shot capabilities, enabling it to extract structures from unknown classes. Our study employed SAM to extract cells from the background in microscopic images containing multiple cells. While SAM demonstrated high accuracy in cases with clear cell contours, its performance was less effective in instances where cell contours were less defined, as shown in Figure 6. It presents challenges in extracting cellular organelles from SBFs with indistinct cell outlines. Therefore, although SAM works well for separating cells from backgrounds, a U-Net model trained to recognize cell organelles from SBF images is needed to overcome its limitations in more complex images.
(a) Large-Scale Microscopic Image
(b) SBF including multi cells
(c) SAM segmentation for Large-Scale Microscopic Image
(d) SAM segmentation for Cross-Sectional Image
(e) Image of sample with background removed
(f) Cell Segmentation Using the 3D Watershed Method.
Figure 6: A schematic diagram of automatically identifying individual cells from a 3D image containing multiple cells using SAM and the 3D Watershed Method. (a), (b) Cross-sectional image of a sample containing multiple cells and its SBF. (c) SAM is used to extract cells from the background. (d) SAM is difficult to classify cellular organelles. (e) Background identified by SAM is removed from the SBF in (b). (f) The 3D watershed method identifies and color-codes individual cells.
3.2 3D watershed algorithm
@@
The next challenge was automating individual cell identification after completing the classification between cells and backgrounds using SAM segmentation. For this task, we employed the 3D watershed method. The Watershed algorithm segments image regions based on brightness values 87344 . This algorithm simulates water flowing from mountains to valleys, dividing the image into different areas. The 3D watershed algorithm extends this concept to three dimensions, enabling the segmentation of the image space by structural region. This method allows for the identification of structures on a per-cell basis. Typically, images are binarized into black and white before implementing this method, clearly distinguishing between areas of high and low brightness values. A specific brightness threshold is needed for this binarization.
@@
Applying the Watershed method directly to noisy images can be challenging, especially when noise obscures the contours of structures, hindering effective segmentation. To address this, we initially used SAM segmentation for large-scale microscopic images containing multiple cells. This preliminary step facilitated image binarization by distinguishing between cells and background. While SAM operates on 2D images and cannot fully determine the continuity between structures in consecutive images, it significantly aids initial segmentation. Following SAM analysis, we applied the 3D Watershed algorithm for further segmentation, making SAM a supportive tool in the 3D Watershed process.

### v3

@@ Abstract
Modeling the 3D structures of cells and tissues is crucial in biology. Sequential cross-sectional images from electron microscopy provide high-resolution intracellular structure information. The segmentation of complex cell structures remains a laborious manual task for experts, demanding time and effort. This bottleneck in analyzing biological images requires efficient and automated solutions.
@@
In this study, the deep learning-based automated segmentation of biological images was explored to enable accurate reconstruction of the 3D structures of cells and organelles. An analysis system for the cell images of Cyanidioschyzon merolae, a primitive unicellular red algae, was constructed. This system utilizes sequential cross-sectional images captured by a focused ion beam scanning electron microscope (FIB-SEM). A U-Net was adopted and training was performed to identify and segment cell organelles from single-cell images. In addition, the segment anything model (SAM) and 3D watershed algorithm were employed to extract individual 3D images of each cell from large-scale microscope images containing numerous cells. Finally, the trained U-Net was applied to segment each structure within these 3D images. Through this procedure, the creation of 3D cell models could be fully automated. The adoption of other deep learning techniques and combinations of image processing methods will also be explored to enhance the segmentation accuracy further.
keywords:
bioimage analysis, segmentation, organelle, cell division, FIB-SEM
Introduction
@@
Advanced deep learning technology has significantly improved the accuracy and efficiency of general image recognition tasks. Furthermore, its application to biological tissue and medical imaging has attracted much attention. As a demonstration, we analyzed 3D cell images captured using a focused ion beam scanning electron microscope (FIB-SEM). FIB-SEMs are primarily used to measure physical properties and evaluate microstructures of solid samples [1]. Recently, they have also been utilized to analyze 3D structures of cells and tissues. The surface of resin-embedded biological samples is repeatedly cut using a gallium ion beam and imaged by a scanning electron microscope. The serial block-face (SBF) image data, i.e., a sequence of cross-sectional images, is reconstructed as a 3D structure.
@@
However, this technology still requires improvement. The intricate boundaries separating structures within each slice, low contrast, and significant noise make analysis based on brightness challenging. Furthermore, as images taken with an electron microscope are single channel (grayscale), relying on electron beam reflections, the absence of color information further complicates the analysis. In addition, identifying specific cells within a space containing multiple cells and extracting their microstructures is difficult. Consequently, the segmentation of microstructure depends on subjective evaluation and manual techniques performed by experts, consuming a considerable amount of time. Image data analysis, which automates this process and is more efficient, objective, and quantitative, is expected to contribute to the analysis of various 3D biological images. The objective of this study was to automate this process fully, exploring the application of deep learning in the recognition tasks of cultured cell images.
@@
We chose Cyanidioschyzon merolae as a model for the automated reconstruction of 3D cell structures using an FIB-SEM [2, 3]. C. merolae is a primitive unicellular eukaryote belonging to red algae, with a diameter of approximately 1.5–2.0 $\mu$m [4] (Figure 1). This organism has a simple structure with only one organelle of each type: nucleus, mitochondrion, plastid, and peroxisome [5]. Each organelle is distributed equally to the two daughter cells [6, 7] (Figure 2). The genome of C. merolae and the genomes of its mitochondrion and plastid have been completely sequenced [8, 9, 10, 11]. Moreover, it is relatively easy to synchronize the cell division cycles of multiple C. merolae cells in culture with the cycle of light [12]. Due to these characteristics, C. merolae is regarded as a model organism for investigating the mechanisms of cell division. Cell division in C. merolae has been extensively studied using molecular and cell biological approaches. Recent advances in observation techniques have enabled the observation of dynamic changes in the behavior and target molecules of individual living cells [13]. However, the physical mechanism of division control based on mechanical properties remains unclear.
@@
Therefore, this article proposes an automated method for accurately extracting cell and organelle structures during the division process of C. merolae. In previous studies, a technique for segmenting each structure from C. merolae FIB-SEM images using machine learning was proposed [14]. However, this method segments structures from each cross-sectional image in the SBF image containing only one cell. Separating cells and backgrounds is still performed manually and is time-consuming. Additionally, the method disregards the sequential contours between cross-section images and has problems in the segmentation of structures at the cell edge. This article proposes a technique to resolve such shortcomings. This method, which enables the automated and rapid acquisition of 3D structures of multiple C. merolae cells, will facilitate the investigation of the mechanical characteristics related to cell division in the future.
Materials and Methods
Image Acquisition and Preprocessing
Image data acquisition using FIB-SEM
@@
The images of freeze-fixed samples containing multiple C. merolae cells were obtained by an FIB-SEM. The aspect ratios of the obtained sequential sectional images were corrected, and the noise was removed using a median filter using Amira software (Thermo Fisher Scientific Inc.). Experts performed cropping to construct learning data images, to create SBF images each containing one cell with side lengths of approximately 200-500 pixels, as depicted in Figure 1 (a). The experts manually segmented each sequential cross-sectional image constituting the SBF image to label the plastid, mitochondrion, peroxisome, nuclei, cytoplasm, and background [14]. In other words, all voxels constituting the SBF images were labeled into six classes.
@@
The cell cycle of C. merolae is classified into five stages [12], shown in Figure 2. The G1-S phase of the cell cycle is referred to as Stage 1, followed by the G2 phase to the M phase (pre-phase) as Stage 2, the M phase (pre-phase) to the M phase (post-phase) as Stage 3, the M phase (post-phase) as Stage 4, and the M phase (termination phase) as Stage 5. Multiple SBF images were labeled for each stage. The labeled SBF images totaled 68 and were classified into one of these stages by experts.
Image data preprocessing
@@
The images obtained by FIB-SEM and the labeled dataset created by experts were preprocessed for deep learning analysis; due to a difference in resolution between the horizontal and vertical directions when capturing images with the FIB-SEM, the voxels in the SBF images were shaped as rectangular parallelepipeds rather than cubes. The SBF images were resized using nearest-neighbor interpolation to restore the aspect ratios of the voxels to their actual size ratios, facilitating 3D data analysis. A bilateral filter technique was employed to process the images and remove noise that Amira software could not eliminate [15]. Each sectional image was resized to create cubic voxels, each side measuring 256 pixels, to ensure a consistent input image size for the deep learning model. For effective deep learning, the brightness values of the electron microscope images were normalized to a range of 0–1 (by dividing 8-bit values by 255). After these preprocessing steps, the dataset was split into different sets for cross-validation, resulting in six cases. In each case, an SBF image was selected as validation data to train the deep learning model, each corresponding to a different stage of cell division. The remaining SBF images were used as training data for the deep-learning model. In addition to the conventional cross-sectional images obtained from the top surface by FIB shaving, the SBF training image dataset was augmented by creating new cross-sectional images from both lateral and anterior views. Details on the number of cells used and the number of cross-sectional images for each dataset in each cross-validation case are presented in Table 1.
Automated Segmentation of Single Cells Using Deep Learning
Deep neural network architecture
@@
We developed a scanning 3D U-Net neural network model, as illustrated in Figure 3 (parameters are shown in Table 2) [16, 17, 18]. Initially created for image segmentation tasks, U-Net is now extensively employed in various fields, including medical imaging [19]. The first half of the neural network consists entirely of convolutional layers. On the opposite side, deconvolutional layers gradually restore the image size to the original input. The convolutional and deconvolutional layers are connected by skip connections, designed to prevent information loss due to the encoding of the convolutional layers. In this study, we implemented four stages of convolution and deconvolution, with each deconvolution layer in our model enhanced with an attention gate mechanism [20]. The ReLU function was used as the activation function at each layer, and batch normalization was performed at all layers to prevent overfitting [16, 17]. The U-Net model in this study had approximately 130 million parameters. We performed a six-class segmentation task. The input data shape was defined as a tensor of [batch size, number of channels, vertical pixel count of the SBF image, and horizontal pixel count of the SBF image]. The label data and U-Net output values were fed into the TverskyLoss function for loss score calculation [21]. The AdamW learning function was used, with the learning rate starting from a base rate of 0.00001 and decaying by a factor of 0.95 each epoch [22]. The number of epochs was set to 5. All networks were built on a GPU (NVIDIA GeForce RTX 4090 with 24 GB memory), and the framework was implemented using Python and TensorFlow. To accelerate learning, the tensor cores of the GPU were automatically utilized.
Segmentation accuracy
@@
During the training phase, we evaluated the accuracy of the learning model through cross-validation. Our cross-validation process included six distinct cases (Table 1); for each case, a dataset of 63 SBF images was used for training, and a separate set of 5 SBF images (one for each stage in the cell cycle) was employed for validation. The SBF images for validation did not overlap between cases. Each validation set included one cell from each of the five stages of the cell division cycle. We segmented the sequential cross-sectional images into six classes using the U-Net for each case. The output of the U-Net approached 1 if the segmentation results indicated a specific class and approached 0 otherwise. Finally, we compared these inference results with the label data and evaluated the model using the intersection over union (IoU) as expressed:
$IoU=\frac{\text{Area of overlap}}{\text{Area of union}}.$ (1)
The area of overlap is the intersection of the predicted and ground truth areas, and the area of union refers to the union of the predicted and ground truth sets.
Automated Extraction of Individual Cells from Large-Scale Images
In this study, we first implemented U-Net-based automated segmentation of organelles using single-cell SBF images that were manually annotated by experts. However, because a model trained solely on single-cell data may have challenges accurately delineating boundaries in large-scale images containing multiple cells, a complementary approach for extracting individual cells from such images is required.
Segment anything model
@@
To generate SBFs automatically, we initially utilized the segment anything model (SAM) to distinguish cells and backgrounds automatically. Developed by Meta, the SAM enables easy segmentation of cells and backgrounds from large-scale microscopic images containing multiple cells [23] (Although SAM2 has recently been announced, this study utilizes SAM, as that was the version available at the time of the research). SAM was trained on over 110 million images using over one billion masks and possesses zero-shot capabilities, enabling it to extract structures from unknown classes. We employed SAM to extract cells from the background in microscopic images containing multiple cells. Although SAM demonstrates high accuracy in cases with clear cell contours, it is less effective in instances in which cell contours are less defined, as shown in Figure 6. This situation presents challenges in extracting cellular organelles from SBFs with indistinct cell outlines. Therefore, although SAM works well for separating cells from backgrounds, a U-Net model trained to recognize cell organelles from SBF images is needed to overcome its limitations in more complex images.
Three-dimensional watershed algorithm
@@
The next challenge was automating individual cell identification after completing the classification between cells and backgrounds using SAM segmentation. For this task, we employed the 3D watershed method. The watershed algorithm segments image regions based on brightness values [24]. This algorithm simulates water flowing from mountains to valleys, dividing the image into different areas. The 3D watershed algorithm extends this concept to three dimensions, enabling the segmentation of the image space by structural region. This method enables the identification of structures on a per-cell basis. Typically, images are binarized into black and white before implementing this method, clearly distinguishing between areas of high and low brightness values. A specific brightness threshold is needed for this binarization.
@@
Applying the watershed method directly to noisy images can be challenging, especially when noise obscures the contours of structures, hindering effective segmentation. To address this issue, we initially performed SAM segmentation for large-scale microscopic images containing multiple cells. This preliminary step facilitated image binarization by distinguishing between cells and background. Although SAM operates on 2D images and cannot fully determine the continuity between structures in consecutive images, it significantly aids initial segmentation. Following SAM analysis, we applied the 3D watershed algorithm for further segmentation, making SAM a supportive tool in the 3D watershed process.
Results
Performance of U-Net Segmentation on SBF Images
Table 3 presents the IoU score for each case based on the inference results of the U-Net. The table organizes six training classes by row and five cell division stages by column. These results indicate that the average accuracy of background identification is approximately 0.99, suggesting that cells and backgrounds can be identified with very high accuracy, as illustrated in Figure 4. Additionally, the accuracy of plastid inference remains high throughout all stages of cell division. In contrast, the mitochondrion and nuclei are smaller in volume than the plastid, making them more difficult to identify accurately, with an average accuracy of about 0.70. The average accuracy for peroxisomes is notably lower, approximately 0.45. Furthermore, the inference accuracy for all classes tends to decrease gradually as the cell division stage progresses.
Plastids are relatively easily identifiable due to their size, unique structure, and distinct brightness values. In contrast, mitochondria and nuclei have brightness values similar to those of the background and starch, making them more challenging to identify. Additionally, peroxisomes are assumed to be difficult for deep learning models to learn due to their irregular shapes that vary from cell to cell and brightness values that are similar to those of the background and starch [25, 26]. Furthermore, as cell division progresses, especially before division, a single SBF image may encompass two cells, increasing the complexity of cell structures within SBF images and potentially decreasing the inference accuracy as cell division advances.
Multi-Directional Inference and Its Impact
We segmented various cellular organelles from cross-sectional images of 2D pictures. However, a problem was identified with low identification accuracy in some classes. To address this issue, we generated new cross-sectional images perpendicular to the original images, using the 3D SBF data, as shown in Figure 5. We used these images for inference with a trained model. This model calculates the probability of belonging to all classes for each voxel (the smallest unit of a 3D image). Three predictive values are obtained for each voxel with inference from three directions. The class with the highest probability is selected when different classes are predicted. As this process may introduce noise, a median filter removes noise after selection. Consequently, as shown in Table 3, some cases exhibit no change or a slight decrease in predictive accuracy, but other samples show clear improvements in accuracy. This finding suggests that accurately determining the class from a cross-sectional image from a specific direction may not be possible. Also, considering that the orientation of the cell to be imaged cannot be predetermined and that inference from difficult angles is inevitable, inferring from three directions is worthwhile, even if it means accepting a slight decrease in accuracy.
Application to Large-Scale Microscopic Image Segmentation